# Cross-Lingual Consistency of Factual Knowledge in Multilingual Language Models

### v1

@@ 1 Introduction
In this work, we take a different stance: instead of measuring the amount of factual knowledge encoded in each language, we focus on its consistency across languages, which shows how knowledge generalizes among different languages in the PLM. As illustrated in Figure 1, the multilingual BLOOM-3b model (Scao et al., 2022) outputs consistently correct completions of the first prompt when queried in English, Spanish, and Vietnamese, but not in Hungarian and Greek. The model also outputs consistent, though wrong, answers to the second query in English, Spanish, and Vietnamese (but not in Hungarian and Greek), suggesting the first three languages share relevant knowledge representations within the model.
@@
We conduct the first in-depth study of CLC of factual knowledge in multilingual PLMs and make the following contributions: (i) We propose a novel Ranking-based Consistency (RankC) metric which assesses knowledge consistency independently from accuracy. (ii) We filter the existing unbalanced datasets (Jiang et al., 2020; Kassner et al., 2021) to form a multi-parallel CLC benchmark, Balanced Multilingual LAnguage Model Analysis (BMLAMA), which has the same set of prompts translated into all languages. (iii) We apply the new metric to BMLAMA to assess CLC in various encoder-only, decoder-only, and encoder-decoder PLMs, including XLM-RoBERTa-large, mT5-large, and BLOOM series. We analyze a number of language properties that correlate with CLC and provide new insights into how factual knowledge percolates among languages. Finally (iv) we present a case study with a state-of-the-art model editing technique based on neuron interpretability (Meng et al., 2022), which provides preliminary evidence that CLC is predictive of whether a fact inserted in language X will transfer to language Y.
@@ 2 Related Work
Since first proposed by LAMA (Petroni et al., 2019), prompt-based probing has become the main technique to assess factual knowledge in PLMs (Davison et al., 2019; Bouraoui et al., 2020; Shin et al., 2020; Brown et al., 2020; Alghanmi et al., 2021; Peng et al., 2022). Given the knowledge represented in a tuple (subject, relation, object), a query $q$ is formed by filling the subject into a relation-specific template, which is fed into the PLM. If the prediction is in line with the object, the model is considered to possess this knowledge. For instance, when queried with ‘The capital of People’s Republic of China is’, the PLM is considered to capture this piece of knowledge if it manages to output the correct answer ‘Beijing’.
@@
Next to a multitude of work focusing on English, a few notable studies have probed factual knowledge multilingually by translating the English prompt-object pairs into a number of languages. X-FACTR (Jiang et al., 2020) and MLAMA (Kassner et al., 2021) indicate strong variations between the amount of knowledge in different languages, due to the size of their training corpora. Apart from English and a handful of other high-resource European languages, very low (i.e. $<$10%) probing accuracies are reported overall. Another relevant work, GeoMLAMA (Yin et al., 2022), probes specifically commonsense knowledge that is susceptible to vary across different regions, leading to the rather surprising finding that the best language to probe knowledge about a certain country (e.g. China) is often not the native language of the given country (e.g. Chinese). The main focus of all these studies is on assessing the amount of factual knowledge encoded for each language, rather than on understanding how such knowledge percolates among languages.
@@
Self-consistency refers to a PLM’s ability to output the same answer to meaning-preserved paraphrases of the same query. Self-consistency of English PLMs has received attention across different tasks (Li et al., 2019; Mitchell et al., 2022; Wang et al., 2022). Fierro and Søgaard (2022) extend the study of self-consistency to multilingual PLMs, by measuring it separately in each language. Their results show poor self-consistency in all languages.
@@
To our knowledge, we are the first to conduct a systematic analysis of cross-lingual consistency of factual knowledge in multilingual PLMs, that is the extent to which a PLM returns the same answer to the same question asked in different languages. As part of their probing study, Jiang et al. (2020) compute the ratio of correct predictions that overlap between two languages within mBERT (cf. Sect. 3.1). They report low ratios overall, with a peak of only 34% in the most similar pair (English-Dutch), but do not further investigate the factors that determine consistency. Moreover, they limit this analysis to one (encoder-only) model while we also examine an encoder-decoder and a series of decoder-only models (cf. Sect. 5.1). As another important difference, we take a more holistic view of consistency, whereby predictions that are incorrect but refer to the same entity across languages should also be considered consistent. Interestingly, concurrent work by Ohmer et al. (2023) proposes to use cross-lingual consistency of a model’s predictions as a means to assess its understanding of meaning beyond specific word forms. They showcase their approach in two language understanding tasks (paraphrase identification and natural language inference). Despite the different scope, their evaluation of ChatGPT using English, German, and Chinese translations reveals limited consistency in the model responses, which is in line with our factual probing findings (cf. Sect. 5) and further indicates this problem persists in very large-scale, last-generation PLMs.
@@ 3 Measuring Cross-Lingual Consistency
Each language $l\in\mathcal{L}$ has a set of queries (i.e. prompts) defined as $Q_{l}$. For each query $q_{i}\in Q_{l}$, there are $N_{i}$ corresponding candidates (Kassner et al., 2021; Yin et al., 2022). For example, the query ‘Steve Jobs worked for __’ has 10 candidates: Apple, Nintendo, Google, WWE, Alexandria, Germany, Yahoo, Berlin, BBC, Microsoft. Each query is fed to the PLM, and the returned probabilities are used to compute a ranking score for each of the candidate words. The specific score calculation depends on the type of model (encoder-only, encoder-decoder, or decoder-only) and on the way a candidate word gets segmented into subwords (see details in Appendix A). After being sorted by ranking scores, the candidate set for $q_{i}$ is represented as $\{c_{i}^{1},\ldots,c_{i}^{N_{i}}\}$, where $c_{i}^{1}$ has the highest prediction probability and $c_{i}^{N_{i}}$ has the lowest. Note that the existing multilingual datasets for knowledge probing (X-FACTR (Jiang et al., 2020) and MLAMA (Kassner et al., 2021)) have different numbers of queries in different languages, which is problematic for measuring consistency.
@@ 3.2 This Work: RankC Metric
where $N_{i}$ is the number of candidates for queries $q_{i}$ and $q^{\prime}_{i}$. [^4] For alternative ranking-based weighting schemes, see Appendix B. Combining equations 3, 4 and 5, the RankC metric becomes:
@@
A computation example is given in Appendix C.
@@
We performed an empirical comparison of RankC to the previously used metric (COverlap, cf. Eq. 1) on the same dataset. The results in Appendix E show that nearly all language pairs with a high COverlap score also receive a high RankC score. Additionally, RankC reveals some new high-consistency pairs that had a low COverlap score due to low probing accuracy.
@@ 4 Experimental Setup
As explained in Sect. 3.2, RankC requires that the queries and their candidates are translated into all evaluated languages. To this end, we extract from X-FACTR (Jiang et al., 2020) and MLAMA (Kassner et al., 2021) all queries that fulfill this criterion. We call the resulting multi-parallel dataset Balanced Multilingual LAnguage Model Analysis (BMLAMA), and release it in two versions: BMLAMA-17 including 6.7k queries in 17 languages (close to X-FACTR which includes 23 languages), and BMLAMA-53 including 3k queries in 53 languages (same as MLAMA). The detailed statistics are shown in Table 1.
@@ 5 Main Consistency Results
Before looking at consistency, we present in Figure 2 the factual probing accuracy results of the three PLMs on BMLAMA-17.555The experiments in this section use BMLAMA-17 as it enables better visualization. Results for BMLAMA-53 are given in Appendix G and discussed in the analysis of Section 6. We first notice that the encoder-only XLM-RoBERTa-large and encoder-decoder mT5-large models outperform the whole decoder-only BLOOM series in terms of average probing accuracy. The trend across languages is similar among the three models, however, BLOOM stands out with an English accuracy that is much higher than all other languages. Regarding model size (BLOOM series, green bars), we find that increasing the number of parameters leads to slight but consistent improvements in factual probing accuracy, which is in line with previous work (Petroni et al., 2019).
Our XLM-RoBERTa-large results are consistent with those reported by Jiang et al. (2020) on X-FACTR, which demonstrates the reliability of our proposed dataset BMLAMA.
@@ 5.1 Consistency in Different PLMs
Figure 3 presents RankC results for the three PLMs. The first observation is that average consistency666 Since the analysis focuses on cross-lingual consistency, we exclude the pairs of one language and itself (e.g. en-en, fa-fa), where RankC values are constantly to be 1. is rather low in all models, and lowest in BLOOM-3b (25%). This negative result is in line with the low overlap ratio of correct predictions observed by Jiang et al. (2020) on mBERT.
@@ 5.2 Effect of Model Size
As observed above (green bars in Figure 2) and by previous work (Petroni et al., 2019), the ability to retrieve correct knowledge grows with model size when other factors are fixed. We ask whether this is also the case for CLC, however the BLOOM results in Figure 4 show only a minor variation in average RankC (+2%) when moving from the smallest to the largest model in our series, i.e. a 5-fold increase in parameters. While this pattern cannot be safely generalized to other models, it does suggest that cross-lingual consistency may remain a problem in very large-scale PLMs.
@@ 6.1 Typological Similarity
Results are shown in Table 2 for both BMLAMA-17 and the smaller but more multilingual BMLAMA-53.999RankC and genetic similarity values on BMLAMA-53 are given in Appendix G. For BMLAMA-17, we find that RankC has a moderate correlation with genetic similarity, a weak correlation with geographical similarity, but no significant correlation with syntactic similarity. As expected, no correlation is observed with phonological similarity. The correlation results on the more comprehensive dataset BMLAMA-53 are similar except for syntactic similarity which acquires a weak positive correlation. Somewhat surprisingly, genetic and geographical similarities see their correlations slightly decreased in this larger dataset, which might be due to the presence of noise in the typological vectors of low-resource languages.
@@ 6.2 Subword Vocabulary Overlap
We consider two corpora: BMLAMA itself and Flores-200 (Costa-jussà et al., 2022). The former is expected to be very relevant, but the resulting correlation may be less generalizable as it is the same corpus on which consistency itself is measured. By contrast, the latter is a mix-domain set of 2k sentences translated from English into 200 languages for the purpose of MT evaluation. Because we are interested in the extent to which different languages make use of the exact same word representations, we segment the corpora with the model’s tokenizer before measuring vocabulary overlap, which makes this metric model-dependent. [^10] See Appendix G for the vocabulary overlap scores of all language pairs in BMLAMA-53.
Shown in Table 2 (right), the Pearson correlation scores on BMLAMA prove that subword vocabulary overlap has a significantly strong impact on the cross-lingual consistency of knowledge in the PLMs, overshadowing the effect of genetic similarity. This suggests that factual knowledge may be primarily percolating across languages in a rather shallow way (through the shared use of some subword embeddings) and conversely, it may be hampered in the absence of such anchors even when languages are related. For instance, the highest-consistency pair in BLOOM-3b is Ukrainian-Russian, which lies close in the language tree (genetic similarity: 0.8) and shares a vast number of subword vocabulary overall (vocabulary overlap: 0.76). However, when querying the working place of David Cameron, BLOOM-3b predicts the correct answer in the Russian query (‘London’) but a wrong answer in Ukrainian (‘Moscow’). This suggests that the correct knowledge did not transfer from Russian to Ukrainian due to the limited subword overlap between these two queries (0.17). When vocabulary overlap is measured on Flores (last column of Table 2), correlation is lower but still significantly strong, indicating our findings are not limited to our benchmark. The correlation between cross-lingual knowledge consistency and vocabulary overlap is clearly illustrated in Figure 5.
@@
Figure 5: Linear regression between subword vocabulary overlap (Flores) and CLC measured on BMLAMA-53 for BLOOM-3b. For results on mT5-large and XLM-RoBERTa-large, see Appendix G.
@@ 7 Case Study: Cross-Lingual Consistency and Knowledge Incorporation
Counterfactual knowledge
Case study
We examine BLOOM-3b since ROME is currently only applicable to decoder-only models. English is chosen as the source language in which the fact is inserted. As target languages, we choose two that have high consistency (RankC) with English (Spanish and Vietnamese) and two with low RankC (Hungarian and Greek). These languages are also diverse in terms of script and relatedness to English. Six queries are picked by ensuring that the PLM selects the originally correct answer as highly probable before editing. Besides, for each edited knowledge, the targeted named entities are the same tokens across all languages. For instance, after inserting ‘Microsoft’ as the workplace of Steve Jobs in English, the output probabilities of the exact tokens, ‘Apple’ and ‘Microsoft’, are calculated when probing with non-English queries. For the evaluation, we follow the setup in Meng et al. (2022) and narrow down candidate sets into two words – one correct and one wrong. The latter is the editing target of ROME. Fed with each query, the PLM calculates logit values for the correct and wrong answers, namely ${\rm logit}_{C}$ and ${\rm logit}_{W}$ respectively. These logits vary vastly among different languages. To focus on the relation between the original and edited fact, we normalize the logits following the previous work (Sarti et al., 2023) as $\frac{{\rm logit}_{C}}{{\rm logit}_{C}+{\rm logit}_{W}}$ and $\frac{{\rm logit}_{W}}{{\rm logit}_{C}+{\rm logit}_{W}}$.  [^11] The raw probability values are provided in Appendix J.
@@
Table 3 shows the results for three queries. A very clear pattern emerges: when a fact is inserted in English, it propagates consistently to the high-CLC languages (i.e. Spanish and Vietnamese). Conversely, the low-CLC languages (Hungarian and Greek) are much less affected and still output higher probabilities for the correct answers even after model editing. The three remaining queries given in Appendix I show the same pattern.

### v5

@@ 1 Introduction
In this work, we take a different stance: instead of measuring the amount of factual knowledge encoded by a PLM in each language, we focus on its consistency across languages. As illustrated in Figure 1, the multilingual BLOOM-3b model (Scao et al., 2022) outputs consistently correct completions of the first prompt when queried in English, Spanish, and Vietnamese, but not in Hungarian and Greek. The model also outputs consistent, though wrong, answers to the second query in English, Spanish, and Vietnamese (but not in Hungarian and Greek), suggesting the first three languages share relevant knowledge representations within the model.
@@
We conduct the first in-depth study of CLC of factual knowledge in multilingual PLMs and make the following contributions: (i) We propose a novel Ranking-based Consistency (RankC) metric which assesses knowledge consistency independently from accuracy. (ii) We filter the existing unbalanced datasets (Jiang et al., 2020; Kassner et al., 2021) to form a multi-parallel CLC benchmark, Balanced Multilingual LAnguage Model Analysis (BMLAMA), which has the same set of prompts translated into all languages. (iii) We apply the new metric to BMLAMA to assess CLC in various encoder-only, decoder-only, and encoder-decoder PLMs, including XLM-RoBERTa-large, mT5-large, and BLOOM series. We analyze a number of language properties that correlate with CLC and provide new insights into how factual knowledge percolates among languages. Finally (iv) we present a case study with a state-of-the-art model editing technique based on neuron interpretability (Meng et al., 2022), providing preliminary evidence that CLC is predictive of whether a fact inserted in language X will transfer to language Y.
@@ 2 Related Work
Since first proposed by LAMA (Petroni et al., 2019), prompt-based probing has become the main technique to assess factual knowledge in PLMs (Davison et al., 2019; Bouraoui et al., 2020; Shin et al., 2020; Brown et al., 2020; Alghanmi et al., 2021; Peng et al., 2022). Given the knowledge represented in a tuple (subject, relation, object), a query $q$ is formed by filling the subject into a relation-specific template, which is fed into the PLM. If the prediction is in line with the object, the model is considered to possess this knowledge. For instance, given a candidate set of city names, when queried with ‘The capital of People’s Republic of China is _’, the PLM is considered to capture this piece of knowledge if it gives the highest probability to the correct answer ‘Beijing’, among all candidates.
@@
Next to a multitude of works focusing on English, a few notable studies have probed factual knowledge multilingually by translating the English prompt-object pairs into a number of languages. X-FACTR (Jiang et al., 2020) and MLAMA (Kassner et al., 2021) indicate strong variations between the amount of knowledge in different languages, due to the size of their training corpora. Apart from English and a handful of other high-resource European languages, very low (i.e. $<$10%) probing accuracies are reported overall. Another relevant work, GeoMLAMA (Yin et al., 2022), probes specifically commonsense knowledge that is susceptible to vary across different regions, leading to the rather surprising finding that the best language to probe knowledge about a certain country (e.g. China) is often not the native language of the given country (e.g. Chinese). The main focus of all these studies is on assessing the amount of factual knowledge encoded for each language, rather than on understanding how such knowledge percolates among languages.
@@
Self-consistency refers to a PLM’s ability to output the same answer to meaning-preserved paraphrases of the same query. Self-consistency of English PLMs has received attention across different tasks (Li et al., 2019; Mitchell et al., 2022; Wang et al., 2023). Fierro and Søgaard (2022) extend the study of self-consistency to multilingual PLMs, by measuring it separately in each language. Their results show poor self-consistency in all languages.
@@
To our knowledge, we are the first to conduct a systematic analysis of cross-lingual consistency of factual knowledge in multilingual PLMs, that is the extent to which a PLM returns the same answer to the same question asked in different languages. As part of their probing study, Jiang et al. (2020) compute the ratio of correct predictions that overlap between two languages within mBERT (cf. Sect. 3.1). They report low ratios overall, with a peak of only 34% in the most similar pair (English-Dutch), but do not further investigate the factors that determine consistency. Moreover, they limit this analysis to one (encoder-only) model while we also examine an encoder-decoder and a series of decoder-only models (cf. Sect. 5.1). As another difference, we take a more holistic view of consistency, whereby predictions that are incorrect but refer to the same entity across languages should also be considered consistent. Interestingly, concurrent work by Ohmer et al. (2023) proposes to use the cross-lingual consistency of a model’s predictions as a means to assess its understanding of meaning beyond specific word forms. They showcase their approach in two language understanding tasks (paraphrase identification and natural language inference). Despite the different scopes, their evaluation of ChatGPT using English, German, and Chinese translations reveals limited consistency in the model responses, which is in line with our factual probing findings (cf. Sect. 5) and further indicates this problem persists in very large-scale, last-generation PLMs.
@@ 3 Measuring Cross-Lingual Consistency
Each language $l\in\mathcal{L}$ has a set of queries (i.e. prompts) defined as $Q_{l}$. For each query $q_{i}\in Q_{l}$, there are $N_{i}$ corresponding candidates (Kassner et al., 2021; Yin et al., 2022). For example, the query ‘Steve Jobs worked for __’ has 10 candidates: Apple, Nintendo, Google, WWE, Alexandria, Germany, Yahoo, Berlin, BBC, Microsoft. Each query is fed to the PLM, and the returned probabilities are used to compute a ranking score for each of the candidate words. The specific score calculation depends on the type of model (encoder-only, encoder-decoder, or decoder-only) and on the way a candidate word gets segmented into subwords (see details in Appendix B). After being sorted by ranking scores, the candidate set for $q_{i}$ is represented as $\{c_{i}^{1},\ldots,c_{i}^{N_{i}}\}$, where $c_{i}^{1}$ has the highest prediction probability and $c_{i}^{N_{i}}$ has the lowest. Note that the existing multilingual datasets for knowledge probing (X-FACTR (Jiang et al., 2020) and MLAMA (Kassner et al., 2021)) have different numbers of queries in different languages, which is problematic for measuring consistency.
@@ 3.2 This Work: RankC Metric
where $N_{i}$ is the number of candidates for queries $q_{i}$ and $q^{\prime}_{i}$. [^4] For alternative ranking-based weighting schemes, see Appendix C. Combining equations 3, 4 and 5, the RankC metric becomes:
@@
A RankC computation example is given in Appendix D and the interpretation of high/low RankC scores is discussed in Appendix E.
@@
We performed an empirical comparison of RankC to the previously used metric (COverlap, cf. Eq. 1) on the same dataset. The results in Appendix F show that nearly all language pairs with a high COverlap score also receive a high RankC score. Additionally, RankC reveals some new high-consistency pairs that had a low COverlap score due to low probing accuracy.
@@ 4 Experimental Setup
As explained in Sect. 3.2, RankC requires that the queries and their candidates are translated into all evaluated languages. We therefore extract from X-FACTR (Jiang et al., 2020) and MLAMA (Kassner et al., 2021) all queries that fulfill this criterion. We call the resulting multi-parallel dataset Balanced Multilingual LAnguage Model Analysis (BMLAMA), and release it in two versions: BMLAMA-17 including 6.7k queries in 17 languages (close to X-FACTR which includes 23 languages), and BMLAMA-53 including 3k queries in 53 languages (same as MLAMA). The detailed statistics are shown in Table 1.
@@ 5 Main Consistency Results
Before looking at consistency, we present in Figure 2 the factual probing accuracy results of the three PLMs on BMLAMA-17.555The experiments in this section use BMLAMA-17 as it enables better visualization. Results for BMLAMA-53 are given in Appendix H and discussed in the analysis of Section 6. We first notice that the encoder-only XLM-RoBERTa-large and encoder-decoder mT5-large models outperform the whole decoder-only BLOOM series in terms of average probing accuracy. The trend across languages is similar among the three models, however, BLOOM stands out with an English accuracy that is much higher than all other languages. Regarding model size (BLOOM series, green bars), we find that increasing the number of parameters leads to slight but consistent improvements in factual probing accuracy, which is in line with previous work (Petroni et al., 2019).
Our XLM-RoBERTa-large results are consistent with those reported by Jiang et al. (2020) on X-FACTR, which demonstrates the reliability of our multi-parallel dataset BMLAMA.
@@ 5.1 Consistency in Different PLMs
Figure 3 presents RankC results for the three PLMs. The first observation is that average consistency [^6] Pairs of a language to itself (e.g. en-en) have always a RankC of 100% and are excluded from the average. is rather low in all models, and lowest in BLOOM-3b (25%). This negative result is in line with the low overlap ratio of correct predictions observed by Jiang et al. (2020) on mBERT.
@@ 5.2 Effect of Model Size
As observed above (green bars in Figure 2) and by previous work (Petroni et al., 2019), the ability to retrieve correct knowledge grows with model size when other factors are fixed. We ask whether this is also the case for CLC. However, the BLOOM results in Figure 4 show only a minor variation in average RankC (+2%) when moving from the smallest to the largest model in our series, i.e. a 5-fold increase in parameters. While this pattern cannot be safely generalized to other models, it does suggest that cross-lingual consistency may remain a problem in very large-scale PLMs. [^8] See Appendix I for additional results on models with 7b parameters.
@@ 6.1 Typological Similarity
Results are shown in Table 2 for both BMLAMA-17 and the smaller but more multilingual BMLAMA-53.101010RankC and genetic similarity values on BMLAMA-53 are given in Appendix H. For BMLAMA-17, we find that RankC has a moderate correlation with genetic similarity, a weak correlation with geographical similarity, but no significant correlation with syntactic similarity. As expected, no correlation is observed with phonological similarity. The correlation results on the more comprehensive dataset BMLAMA-53 are similar except for syntactic similarity which acquires a weak positive correlation. Somewhat surprisingly, genetic and geographical similarities see their correlations slightly decreased in this larger dataset, which might be due to the presence of noise in the typological vectors of low-resource languages.
@@ 6.2 Subword Vocabulary Overlap
We consider two corpora: BMLAMA itself and Flores-200 (Costa-jussà et al., 2022). The former is expected to be very relevant, but the resulting correlation may be less generalizable as it is the same corpus on which consistency itself is measured. By contrast, the latter is a mix-domain set of 2k sentences translated from English into 200 languages for the purpose of MT evaluation. Because we are interested in the extent to which different languages make use of the exact same word representations, we segment the corpora with the model’s tokenizer before measuring vocabulary overlap, which makes this metric model-dependent. [^11] See Appendix H for the vocabulary overlap scores of all language pairs in BMLAMA-53.
Figure 5: Linear regression between subword vocabulary overlap (Flores) and CLC measured on BMLAMA-53 for BLOOM-3b. For results on mT5-large and XLM-RoBERTa-large, see Appendix H.
Shown in Table 2 (right), the Pearson correlation scores on BMLAMA prove that subword vocabulary overlap has a significantly strong impact on the cross-lingual consistency of knowledge in the PLMs, overshadowing the effect of genetic similarity. This suggests that factual knowledge may be primarily percolating across languages in a rather shallow way (through the shared use of some subword embeddings) and conversely, it may be hampered in the absence of such anchors even when languages are related. For instance, the highest-consistency pair in BLOOM-3b is Ukrainian-Russian, which lies close in the language tree (genetic similarity: 0.8) and shares a vast number of subword vocabulary overall (vocabulary overlap: 0.76). However, when querying the working place of David Cameron, BLOOM-3b predicts the correct answer in the Russian query (‘London’) but a wrong answer in Ukrainian (‘Moscow’). This suggests that the correct knowledge did not transfer from Russian to Ukrainian due to the limited subword overlap between these two queries (0.17). When vocabulary overlap is measured on Flores (last column of Table 2), correlation is lower but still significantly positive, indicating our findings are not limited to our benchmark. The correlation between cross-lingual knowledge consistency and vocabulary overlap is clearly illustrated in Figure 5.
@@ 7 Case Study: Cross-Lingual Consistency and Knowledge Incorporation
Counterfactual knowledge
Case study
We examine BLOOM-3b since ROME is currently only applicable to decoder-only models. English is chosen as the source language in which the fact is inserted. As target languages, we choose two that have high consistency (RankC) with English (Spanish and Vietnamese) and two with low RankC (Hungarian and Greek). These languages are also diverse in terms of script and relatedness to English. Six queries are picked by ensuring that the PLM selects the originally correct answer as highly probable before editing. We also ensure that, for each edited knowledge, the subject and object entities are the same tokens across all languages. This eliminates the concern that, e.g., Spanish and Vietnamese get consistent predictions with English simply because of the lexical co-occurrence of their subject and object tokens in the evaluated queries. For the evaluation, we follow the setup in Meng et al. (2022) and narrow down candidate sets into two words – one correct and one wrong. The latter is the editing target of ROME. Fed with each query, the PLM calculates logit values for the correct and wrong answers, namely ${\rm logit}_{C}$ and ${\rm logit}_{W}$ respectively. These logits vary vastly among different languages. To focus on the relation between the original and edited fact, we normalize the logits following the previous work (Sarti et al., 2023) as $\frac{{\rm logit}_{C}}{{\rm logit}_{C}+{\rm logit}_{W}}$ and $\frac{{\rm logit}_{W}}{{\rm logit}_{C}+{\rm logit}_{W}}$.  [^12] The raw probability values are provided in Appendix K.
Table 3 shows the results for three queries. A very clear pattern emerges: when a fact is inserted in English, it propagates consistently to the high-CLC languages (i.e. Spanish and Vietnamese). Conversely, the low-CLC languages (Hungarian and Greek) are much less affected and still output higher probabilities for the correct answers even after model editing. The three remaining queries given in Appendix J show the same pattern.